OpenStudy (anonymous):

What is the GCF for the polynomial -10x3y4 - 16x5y2?

OpenStudy (cwrw238):

the GCF greatest common factor is the highest factor which will divide inot all the terms so for -10 and -16 it is -2 what do you think the GCF of x^3 and x^5 is?

OpenStudy (anonymous):

3 and 5 can't go into each other evenly like 2 and 5 can. So would x^3 and x^5 just be left alone?

OpenStudy (anonymous):

2 and 4 can *** sorry meant to type 4 instead of 5

OpenStudy (cwrw238):

it is easier for letters than it is for numbers x^3 will divide into itself once and it will also divide into x^5 x^5 / x^3 = x^(5-3) = x^2 so GCF of x^3 and x^5 = x^3

OpenStudy (anonymous):

thank you so much, i understand now. :)

OpenStudy (cwrw238):

yw